a right triangle has a leg that is 5 in. long and a hypotenuse that is 13 in long. How do you find the length of the 3rd side

Answer by nyc_function(2741)

Use the Pythagorean Theorem.
Let s = third side length
5^2 + s^2 = 13^2
25 + s^2 = 169
s^2 = 169 - 25
s^2 = 144
sqrt[s^2] = sqrt[144]
s = 12